For the spontaneous as written at 25°C. A) Cu(s) I Cu (0.10 M)IFe (0.0053 M) I Fe(s) O Spontaneous O Not B) Pt(s) ISn2(0.0021 M), Sn (0.15 M) II Fe2 (0.0012 M), Fe (0.12 M) I Pt(s) O Not Spontaneous
Answer #1

A) Cu(s) + Fe2+(aq) ----> Cu2+(aq) + Fe(s)

   Ecell = E0cell-0.0591/nlog[Cu2+/Fe2+]

E0cell = E0cathode - E0anode

        = -0.440 - 0.339

        = -0.779 v

Ecell = (-0.779)-(0.0591/2)log(0.1/0.0053)

       = -0.817 v

DG = - nFEcell

    = -2*96500*-0.817

    = 157.68 kj

as DG = +ve, the process is non-spontaneous.

B) Sn2+ + 2 Fe3+ -----> Sn4+ + 2 Fe2+


Ecell = E0cell-0.0591/nlog ([Fe2+]^2[Sn4+]/[Fe3+]^2[Sn2+])

E0cell = 0.771 - 0.154

         = 0.617 v

Ecell = 0.617-(0.0591/2)log((0.0012^2*0.15)/(0.12^2*0.0021))

       = 0.68 v

DG = - nFEcell

    = -2*96500*0.68

    = -131.24 kj

as DG = -ve, the process is spontaneous.